Circular No. 8525

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2005 J1 (McNAUGHT) R. H. McNaught reports his discovery of a comet with the 0.5-m Up-

psala Schmidt telescope in the course of the Siding Spring Survey; the discovery images (initial observation provided below) show the object as diffuse, and images taken on May 5.7 UT show it to be slightly diffuse with a faint tail ∼ 10′′ long to the west. Following posting on the ‘NEO Confirmation Page, A. C. Gilmore reports that CCD images taken on May 5.5–5.7 with the Mt. John 1.0-m reflector show the object to be slightly more diffuse than the images of surrounding stars.

2005 UT  α_{2000}  δ_{2000}  Mag.
May 3.72216  20^{h}15^{m}57.^{s}78  −28^{o}50′37′′.9  17.6

The available astrometry, the following preliminary elliptical orbital elements, and an ephemeris appear on MPEC 2005-J32.

T = 2005 Apr. 17.402 TT  ω = 338.^{o}980
e = 0.58018  Ω = 268.928 }2000.0
q = 1.53305 AU  i = 31.824
a = 3.65166 AU n^{o} = 0.141244 P = 7.0 years

V1187 SCORPII D. K. Lynch, R. W. Russell, and R. J. Rudy, Aerospace Corporation,

report that spectroscopy (range 0.8–2.5-μm) of V1187 Sco (cf. IAUC 8380, 8382), taken with the Infrared Telescope Facility (+ SpeX) on Apr. 18.7 UT, shows that the object had faded considerably (K = 13.2 from the spectrum) and that the lines were still broad. H I and He I 1.0830- and 2.0581-μm were present (2900 km/s FWHM), as were some weak He II lines. There was a surprising lack of coronal lines, although the [Si VI] 1.9645-μm line was strong and appeared to be split into two strong, wellseparated components. The unidentified novae lines at 1.19 and 1.55 μm were present. There was no evidence of thermal emission from dust.

COMET C/1997 J5 (SOHO) Another faint Kreutz sungrazer (cf. IAUC 8524), stellar in appearance:

Comet 1997 UT α_{2000} δ_{2000} Inst. F MPEC
C/1997 J5 May 2.197 2^{h}48.^{m}8 +12^{o}03′ C3 RK 2005-H24
2005 May 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green